# MagicDriveDiT: High-Resolution Long Video Generation for Autonomous Driving with Adaptive Control

### v2

@@ 7 Conclusion and Discussion
In this paper, we presented MagicDriveDiT, an innovative framework for high-resolution and long video synthesis with precise control, specifically tailored for applications like autonomous driving. By introducing MVDiT block and spatial-temporal conditional encoding module, MagicDriveDiT effectively addresses the challenges of scalability and geometry control in video generation with the DiT architecture and 3D VAE. Our approach also introduces richer contextual description for driving video generation, and a progressive training strategy with variable length and resolution adaption, enhancing the model’s generation ability. Extensive experiments demonstrate that MagicDriveDiT generates realistic videos that maintain spatial and temporal coherence, significantly surpassing previous methods in resolution and frame count, opening new possibilities for simulations and other applications in autonomous driving.
@@
Please find the videos on our project website: https://flymin.github.io/magicdrivedit/
@@ B More Details for Mixed Resolution and Frames Training
MagicDriveDiT is trained through a progressive training approach with variable length and resolution data configurations (see Section 4.5). Consequently, our method of data mixing corresponds to the three training stages, as detailed in Table I.
@@ C Efficiency of Progressive Bootstrap Training
Table II: Speed for Each Training Stage of MagicDriveDiT, measured on NVIDA A800 GPUs. Over a 4-day period (for example), Stage 1 training yields nearly 60 times more iterations than Stage 3, and Stage 2 offers about 7 times more. This value is calculated by multiplication with sequence parallel (SP) size (in practice, we use SP size of 4 for the stage 3, with 66.24s/it).
@@ F Human Evaluation for Multi-frame & Multi-view Consistency
Evaluating multi-frame and multi-view consistency in video generation has long been a challenging issue [1], as the academic community lacks a unified standard to accurately assess such consistency. To address this, we employed human evaluation to measure these two aspects of consistency. Specifically, we invited participants with diverse backgrounds to compare videos generated by MagicDriveDiT and Gao et al. [11] under identical conditions. Participants were asked to select the video with better consistency, and the winning probability of each model was statistically analyzed. As illustrated in Figure II, MagicDriveDiT demonstrated significantly superior consistency, indicating a substantial improvement over Gao et al. [11].
Figure II: Experts evaluated content consistency against Gao et al. [11], showing MagicDriveDiT generates videos with superior consistency.
@@ G More Comparison among VAEs
Table III: VAE Comparison for Street Views. CogVAE [42] and Open-Sora [46] (1.2) are 3D VAEs; SD VAE [30] is 2D VAE, which is also widely adopted by previous street view generation (e.g., [11]). Results are PSNRs calculated through videos from the nuScenes validation set. MagicDriveDiT adopts CogVAE.
@@ H Reason for Using CogVAE without the Pre-trained Diffusion Model
(b) Generation from MagicDriveDiT. We take conditions as inputs and generate the full video (only show the first 9s for comparison).
Figure III: Comparison between Rollout for Long Videos (Vista [12]) and Single Inference (our MagicDriveDiT). Although rollout can handle long videos, the quality is significantly degraded. In contrast, our extrapolation maintains high quality in long video generation.
To achieve long video generation, previous work typically employs a method of future frame prediction combined with rollout. This involves, after the $n$-th inference, taking the last $l$ frames from this inference as the first $l$ frames for the $n+1$-th inference, thus enabling long video generation. However, since the model does not directly capture long-term dependencies and accumulates errors with each inference, such rollouts often fail to support sufficiently long videos. Among rollout methods, Vista [12] currently achieves relatively good results. We compared a 9-second video generated by performing 4 rollouts with Vista (the paper claims it can support 6 rollouts) to a 9-second segment produced by our method, MagicDriveDiT. It is evident that our method maintains consistent video quality over long sequences, whereas Vista’s results show a noticeable decline. Therefore, we believe that the hybrid training and length extrapolation approach adopted by MagicDriveDiT can achieve higher quality in long video generation.
@@
As said in the “Note”, MagicDriveDiT is capable of generating $6\times 848\times 1600\times 241$ videos (20s at 12 fps). We include more generated samples in Figure IV-V. Please see the videos on our project page.

### v3

@@ 7 Conclusion and Discussion
In this paper, we presented MagicDrive-V2, an innovative framework for high-resolution and long video synthesis with precise control, specifically tailored for applications like autonomous driving. By introducing MVDiT block and spatial-temporal conditional encoding module, MagicDrive-V2 effectively addresses the challenges of scalability and geometry control in video generation with the DiT architecture and 3D VAE. Our approach also introduces richer contextual description for driving video generation, and a progressive training strategy with variable length and resolution adaption, enhancing the model’s generation ability. Extensive experiments demonstrate that MagicDrive-V2 generates realistic videos that maintain spatial and temporal coherence, significantly surpassing previous methods in resolution and frame count, opening new possibilities for simulations and other applications in autonomous driving.
@@
Please find the videos on our project website: https://flymin.github.io/magicdrive-v2/
@@ B More Details for Mixed Resolution and Frames Training
MagicDrive-V2 is trained through a progressive training approach with variable length and resolution data configurations (see Section 4.5). Consequently, our method of data mixing corresponds to the three training stages, as detailed in Table I.
@@ C Efficiency of Progressive Bootstrap Training
Table II: Speed for Each Training Stage of MagicDrive-V2, measured on NVIDA A800 GPUs. Over a 4-day period (for example), Stage 1 training yields nearly 60 times more iterations than Stage 3, and Stage 2 offers about 7 times more. This value is calculated by multiplication with sequence parallel (SP) size (in practice, we use SP size of 4 for the stage 3, with 66.24s/it).
@@ F Human Evaluation for Multi-frame & Multi-view Consistency
Evaluating multi-frame and multi-view consistency in video generation has long been a challenging issue [1], as the academic community lacks a unified standard to accurately assess such consistency. To address this, we employed human evaluation to measure these two aspects of consistency. Specifically, we invited participants with diverse backgrounds to compare videos generated by MagicDrive-V2 and Gao et al. [11] under identical conditions. Participants were asked to select the video with better consistency, and the winning probability of each model was statistically analyzed. As illustrated in Figure II, MagicDrive-V2 demonstrated significantly superior consistency, indicating a substantial improvement over Gao et al. [11].
Figure II: Experts evaluated content consistency against Gao et al. [11], showing MagicDrive-V2 generates videos with superior consistency.
@@ G More Comparison among VAEs
Table III: VAE Comparison for Street Views. CogVAE [42] and Open-Sora [46] (1.2) are 3D VAEs; SD VAE [30] is 2D VAE, which is also widely adopted by previous street view generation (e.g., [11]). Results are PSNRs calculated through videos from the nuScenes validation set. MagicDrive-V2 adopts CogVAE.
@@ H Reason for Using CogVAE without the Pre-trained Diffusion Model
(b) Generation from MagicDrive-V2. We take conditions as inputs and generate the full video (only show the first 9s for comparison).
Figure III: Comparison between Rollout for Long Videos (Vista [12]) and Single Inference (our MagicDrive-V2). Although rollout can handle long videos, the quality is significantly degraded. In contrast, our extrapolation maintains high quality in long video generation.
To achieve long video generation, previous work typically employs a method of future frame prediction combined with rollout. This involves, after the $n$-th inference, taking the last $l$ frames from this inference as the first $l$ frames for the $n+1$-th inference, thus enabling long video generation. However, since the model does not directly capture long-term dependencies and accumulates errors with each inference, such rollouts often fail to support sufficiently long videos. Among rollout methods, Vista [12] currently achieves relatively good results. We compared a 9-second video generated by performing 4 rollouts with Vista (the paper claims it can support 6 rollouts) to a 9-second segment produced by our method, MagicDrive-V2. It is evident that our method maintains consistent video quality over long sequences, whereas Vista’s results show a noticeable decline. Therefore, we believe that the hybrid training and length extrapolation approach adopted by MagicDrive-V2 can achieve higher quality in long video generation.
@@
As said in the “Note”, MagicDrive-V2 is capable of generating $6\times 848\times 1600\times 241$ videos (20s at 12 fps). We include more generated samples in Figure IV-V. Please see the videos on our project page.